### v3

# Tug-of-war between idiom’s figurative and literal meanings in LLMs

Abstract
Idioms present a unique challenge for language models due to their non-compositional figurative meanings, which often strongly diverge from the idiom’s literal interpretation. This duality requires a model to learn representing and deciding between the two meanings to interpret an idiom in a figurative sense, or literally. In this paper, we employ tools from mechanistic interpretability to trace how a large pretrained causal transformer (LLama3.2-1B-base) deals with this ambiguity. We localize three steps of idiom processing: First, the idiom’s figurative meaning is retrieved in early attention and MLP sublayers. We identify specific attention heads which boost the figurative meaning of the idiom while suppressing the idiom’s literal interpretation. The model subsequently represents the figurative representation through an intermediate path. Meanwhile, a parallel bypass route forwards literal interpretation, ensuring that a both reading remain available. Overall, our findings provide a mechanistic evidence for idiom comprehension in an autoregressive transformer. [^1] Code and dataset are available at https://github.com/sori424/idiom_processing under the MIT License.
$\dagger$\$$\dagger$\$footnotetext: Co-corresponding authors.
1 Introduction
Figure 1: The figurative and literal interpretations are highlighted in the block and path. We find three main steps for idiom processing: 1 Idiom retrieval step: Early layers (i.e., 0-3) attention and MLP are actively retrieving the idiom’s figurative meaning while storing both figurative and literal interpretations in the residual stream. 2 Selective interpretation step: At the token immediately following the idiom span, the model begins to encode a representation that favors the figurative interpretation over the literal one, starting from the middle layers. 3 Interpretation routing: For final prediction, the model passes literal interpretation via both a direct compositional semantic path ( 3-2), as well as the intermediate pathway that prioritizes the figurative meaning ( 3-1 figurative path).
Idioms challenge standard semantic composition because they are, by definition, multi-word sequences whose figurative meanings diverge from the composition of their literal parts Beck and Weber (2020). For instance, “kick the bucket” conveys the figurative meaning “to die”, yet its literal sense “physically kicking a container” remains semantically available. This ambiguity often leads large language models (LLMs) to misinterpret idioms, favoring their literal sense over the intended figurative one Kabra et al. (2023); Liu et al. (2023); Knietaite et al. (2024). Psycholinguistics has been investigated how humans understand idioms, whether they access the idiomatic meaning directly or first have to access the literal meaning of the idiom and only in a later stage suppress that meaning and access the figurative meaning. These findings have lead to various models describing human idiom processing Gibbs (1980); Cacciari and Glucksberg (1991); Bobrow and Bell (1973); Cacciari and Tabossi (1988).
Yet for LLMs, it’s unclear how they process idioms, potentially limiting further improvements in their idiom comprehension. One explanation may lie in how the model transforms raw input token embeddings into richer representations Tian et al. (2023); Haviv et al. (2022); Dankers et al. (2022). Within the autoregressive transformer, token representations are incrementally refined from the initial input embeddings to the final representation used for next token prediction Vig et al. (2020). This transformation is achieved through a series of residual layers, which consist of multi-head self-attention (MHSA) and MLP sublayers, each adding their outputs to update the representation. This multi-stage transformation suggests that idiom processing might follow a multi-step transformation process. Haviv et al. (2022) demonstrated that the early layers of transformer models store and retrieve idiomatic information, while the higher layers enhance model confidence. Meanwhile, Dankers et al. (2022) showed that during machine translation of idioms, the transformer model recognizes idiomatic expressions and processes them more as single lexical units. This involves grouping the idioms together using attention sublayers and reducing their interaction with surrounding context. However, both studies lack a detailed mapping of the specific model components and causal pathways underlying these idiom processing behaviors.
In this paper, we first aim to identify how the idiom is interpreted by the transformer (i.e., LLama3.2-1B-base), where its figurative meaning is located, and how it is retrieved by the model. To this end, we pinpoint specific components of the model that have specialized in processing idioms by boosting their figurative meaning, while suppressing the compositional semantic literal interpretation. Having identified these, we proceed to investigating the information flow in the model, with the goal of identifying specific pathways through which the figurative meaning is passed forward to the final prediction, and whether this differs from the pathways of passing forward information about an idiom’s literal interpretation.
We address these questions through knockout analyses Nanda et al. (2023); Wang et al. (2022); Geva et al. (2023), where we separately knock out activations of each component (layer-wise MHSA and MLP, individual attention heads) in the model to observe their importance for retrieving idiom’s figurative meaning. Furthermore, to trace the flow of these interpretations, we employ activation patching experiments Wang et al. (2022); Meng et al. (2022); Hanna et al. (2023); Conmy et al. (2023); Stolfo et al. (2023), in which we ablate one or both interpretations from specific pathways within the model.
Our main findings are summarized in Figure 1: we identify three main steps of idiom processing (1) retrieval of the figurative meaning of the idiom in the early layers, (2) selective encoding of the two interpretations at intermediate layers after the idiom span token position, and (3) two alternative paths for routing the idiom’s meaning to the final prediction, via a direct path for the literal meaning and an intermediate pathway that preferentially amplifies figurative interpretation over the other. In summary, our main contributions are:
  • We propose a next token prediction task that distinguishes figurative vs. literal interpretations using interpretation-specific tokens (Section 2).
  • We show that early layers contribute to retrieve the idiom’s figurative meaning, and that these are significantly different from general semantic processing (Section 4).
  • We identify a selection mechanism at the token immediately following the idiom, where its residual activation shifts toward the figurative interpretation and then propagates through the intermediate layers (Section 5).
  • We also uncover a bypass pathway that channels the literal interpretation directly into the final prediction, circumventing the intermediate route (Section 5.2).
2 Idiom interpretation task
We are interested in teasing apart the language model’s processing of the literal vs. figurative interpretation of ambiguous idiom expressions. We therefore formalize a task and construct a dataset to investigate the interpretation shift in the model. In order to allow for effective probing, we embed the idiom into a sentence template, constructed such that the interpretation is revealed via next token prediction. The template [^2] X is instantiated with various pronouns (he/she/it/they); the word “would” is inserted only in some of the idioms to make the sentence sound more fluent; idioms are morphologically fit into the sentence; depending on sg/pl form of X, the second part of the template contained was or were. ‘X (would) IDIOM because X was so/too/a/the’ can be applied to all of the selected idioms while preserving the idiom’s ambiguity. It elicits a causal completion, which we expect to differ between the literal and the figurative meaning of the idiom: for the figurative meaning of the idiom “kick the bucket”, we would expect completions that provide causes for the figurative meaning “to die”, e.g., old/sick, while for the literal meaning, we would expect completions that provide reasons for “kicking an object”, e.g. angry/mad. Figure 2 illustrates this process and selected examples can be found in Appendix Table 2.
Figure 2: Pipeline for extracting figurative ($C_{f}$) and literal ($C_{l}$) candidate tokens for an idiomatic sentence.
Since there is no single ground truth continuation that reveals either the figurative or the literal reading, we automatically construct sets of multiple tokens that are consistent with either the literal or the figurative meaning. The tokens in each set are mutually exclusive and selected based on their high likelihood of being associated with either the figurative or literal meaning of the idiom. In the following, we describe each step of the process of generating the dataset.
1. Idiom Extraction: We select 245 idioms from a psycholinguistic paper Cronk et al. (1993), each of which received high literality ratings in a human rating study.
2. Sentence Generation: We generate a literal paraphrase and a figurative paraphrase for each idiom using the LLama3.3-70B-instruct model Grattafiori et al. (2024). The sentences below are three examples of the generated ambiguous sentence ($s_{a}$), figurative paraphrase ($s_{f}$), and literal paraphrase ($s_{l}$).
[…]
.[$s_{a}$] He would kick the bucket because he was so
[…]
.[$s_{f}$] He would die because he was so
[…]
.[$s_{l}$] He would kick the container because he was so
3. Token Set Generation: To generate sets of next-word completions that are indicative of the figurative vs. the literal interpretation, we prompt the LLaMa3.3-70B model with disambiguated paraphrases ($s_{f}$, $s_{l}$) and obtain logits from the model with greedy decoding. Let $z^{(f)}$ and $z^{(l)}$ denote the logits corresponding to $s_{f}$ and $s_{l}$ as input, respectively. We then compute the element-wise difference between the two sets of logits as follows:
$\Delta z_{v}=z_{v}^{(f)}-z_{v}^{(l)},$
where $v\in V$ is the vocabulary size of the model (i.e., $V=$128,256). Since tokens that score high in one context tend to score low in the other, $\Delta z_{v}$ captures each token’s context-specific relevance. The 20 tokens that have highest $\Delta z_{k}$ are chosen as the set $C_{f}$ of tokens that indicate a figurative interpretation and the 20 tokens that have lowest $\Delta z_{k}$ values are chosen as the set $C_{l}$ that indicates a literal interpretation.
4. Validation: To make sure that the Llama3.2-1B-base model has prior knowledge of the selected idioms, and the capability of discerning the context of the given sentence, we filter out the generated instances that satisfy the following two conditions:
(1) Given $s_{a}$ as a prompt, the cumulative probabilities of predicting a figurative candidate as a next token must exceed that of the literal candidate. This metric quantifies that the model has the knowledge of the idiom:
$\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{a})>\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{a}).$
(2) Given $s_{f}$ and $s_{l}$ as prompt, the cumulative probabilities of predicting next token candidate matching the context exceeds that of the opposite context:
$\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{f})>\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{f}),$
$\sum_{c_{l}\in C_{l}}p(c_{l}\mid s_{l})>\sum_{c_{f}\in C_{f}}p(c_{f}\mid s_{l}).$
110 out of the original set of 245 idioms passed these criteria. All following analyses are based on this subset of 110 validated idioms.
3 Methodology
3.1 Information flow in transformers
Given an input token sequence $\mathbf{t}=[t_{1},\dots,t_{T}]$ over a vocabulary $V$, each token is embedded as $\mathrm{x}_{i}^{0}=\mathrm{RoPE}(\mathrm{emb}(t_{i}),i)\in\mathbb{R}^{d}$, where $\mathrm{RoPE}$ applies rotary positional encoding to inject position $i$ Su et al. (2024), and $d$ is the model’s hidden dimension. From layer 0, these vectors $\mathrm{x}_{i}^{0}$ are carried forward and accumulated over the subsequent $L$ layers. At each layer $\ell$, the hidden state for token $i$ is updated from its previous value $\mathrm{x}_{i}^{\ell-1}$ by adding the multi-head self-attention (MHSA) output $a_{i}^{\ell}$ and MLP output $m_{i}^{\ell}$:
$\mathrm{x}_{i}^{\ell}=\mathrm{x}_{i}^{\ell-1}+a_{i}^{\ell}+m_{i}^{\ell}.$
MLP SublayersEvery MLP sublayer computes a local update for each representation:
$m_{i}^{\ell}=W_{F}^{\ell}\,\sigma\Bigl{(}W_{I}^{\ell}(a_{i}^{\ell}+\mathrm{x}_{i}^{\ell-1})\Bigr{)},$
where $W_{I}^{\ell}\in\mathbb{R}^{d_{ff}\times d}$ and $W_{F}^{\ell}\in\mathbb{R}^{d\times d_{ff}}$ are parameter matrices with inner-dimension $d_{ff}$ and $\sigma$ is a nonlinear activation function.
MHSA SublayersEach attention layer’s output $\mathrm{a}^{\ell}_{i}$ aggregates information of multiple parallel attention heads $h$. The parameter matrices of each MHSA sublayer comprise three projection matrices $W^{Q},\,W^{K},\,W^{V}\in\mathbb{R}^{d\times d}$, and $W_{O}\in\mathbb{R}^{d\times d}$. The columns of each projection matrix and rows of the outputs matrix can be split into $H$ equal parts, corresponding to the number of attention heads $W_{Q}^{\ell,j},\,W_{K}^{\ell,j},\,W_{V}^{\ell,j}\in\mathbb{R}^{d\times\tfrac{d}{H}}$, and $W_{O}^{\ell,j}\in\mathbb{R}^{\tfrac{d}{H}\times d},\,j=[1,H]$. Let $X^{\ell-1}\in\mathbb{R}^{N\times d}$ be the matrix of token representations at layer $\ell-1$. This allows the MHSA output as a sum of matrices, each induced by single attention head:
$a_{i}^{\ell}=\sum_{j=1}^{H}A^{\ell,j}\bigl{(}X^{\ell-1}\,W_{V}^{\ell,j}\bigr{)}\,W_{O}^{\ell,j}.$
After $L$ layers, the final hidden state $\mathrm{x}_{T}^{L}$ is mapped to vocabulary logits via an unembedding matrix $W_{U}\in\mathbb{R}^{|V|\times d}$, so that
$y=W_{U}\,\mathrm{x}_{T}^{L}\in\mathbb{R}^{|V|}$
and $\mathrm{softmax}(y)$ defines the next-token distribution. Thus we can view a transformer as a computational graph $G\colon\mathcal{X}\to\mathcal{Y}$, where token embeddings, MHSA sublayers, MLP sublayers, and the unembedding all communicate via the residual stream. We refer to Vaswani et al. (2017); Geva et al. (2023) for further details.
3.2 Knockout
If we think of a model as a computational graph $G$, then certain behaviors of the model can be attributed to specific subgraphs $G_{sub}\subseteq G$ Rai et al. (2024). Identifying such subgraphs provides a valuable mechanism for interpretability, enabling the isolation of individual components and their contributions to the model’s behavior. A knockout removes specific components to quantify their contribution to particular behaviors Nanda et al. (2023); Wang et al. (2022); Geva et al. (2023). Every activation in $G_{\text{sub}}$ can be replaced by either zero or reference mean:
$\mathrm{x}_{i}^{\ell}\;\longmapsto\;\tilde{\mathrm{x}}_{i}^{\ell}=\begin{cases}0\quad\text{or}\quad\mu_{i}^{\ell}&\text{if }(\ell,i)\in G_{\text{sub}},\\ \mathrm{x}_{i}^{\ell}&\text{otherwise}.\end{cases},$
then $G$ with those components “knocked out” is
$G_{\text{knock}}=G\left(\dots,\tilde{\mathrm{x}}_{i}^{\ell},\dots\right).$
By replacing the activations of these components, we eliminate the contributions of $G_{sub}$ while keeping other computations in $G$ fixed.
3.3 Activation patching
Activation patching is a technique to identify which activations in a model contribute to a particular output Wang et al. (2022); Meng et al. (2022); Hanna et al. (2023); Conmy et al. (2023); Stolfo et al. (2023). Instead of knocking out a component, we replace it with specific representations. This involves running the model on input A, then selectively replacing (patching) certain activations with those obtained from a paired input B:
$G_{\text{patch}}(A\hookleftarrow B)=G\bigl{(}\mathrm{x}^{1}_{i}(A),\dots,\mathrm{x}^{\ell}_{i}(B),\mathrm{x}^{\ell+1}_{i}(A),\dots\bigr{)}.$
By comparing the original output for input A with the altered output after patching, we can precisely measure how much the introduced activations from input B shift the model’s output.
3.4 Metric for measuring interpretation
To measure whether the model assumes the figurative or literal interpretation when given each prompt $s\in\{s_{a},s_{f},s_{l}\}$, we define two scores based on cumulative probabilities:
$F(s)\;=\;\sum_{c_{f}\in C_{f}}p(c_{f}\mid s);\;L(s)\;=\;\sum_{c_{l}\in C_{l}}p(c_{l}\mid s),$
where $C_{f}$ and $C_{l}$ are the sets of tokens corresponding to figurative and literal interpretation, respectively. After applying an intervention (knockout or activation patching), we quantify its causal effect:
$\Delta I(s)\;=\;I_{\mathrm{interv}}(s)\;-\;I_{\mathrm{origin}}(s),\,I\in\{F,L\},$
where $\Delta I<0$ means the intervention disrupts the corresponding interpretation, showing the targeted component is essential for that meaning. In contrast, $\Delta I=0$ indicates no effect and $\Delta I>0$ implies that the intervened component is not required (or may even been inhibited) that particular reading.
4 Localizing idiom meaning retrieval
4.1 Probing sublayers through knockout
To locate the mechanisms for retrieving figurative meanings from the model’s parameters, we knock out the activations at idiom span tokens by replacing them with their mean values computed over all $s_{a}$. By knocking out the activations across each layer, we assess how the probabilities of predicting specific candidate tokens ($\Delta F(s_{a}),\,\Delta L(s_{a})$) change. To narrow down the points to which we can ascribe a functional role for a specific computation component, we limit the intervention to the MLP and MHSA in each block.
(a) MLP
(b) MHSA
(c) MLP
(d) MHSA
Figure 3: Sublayer-wise interpretation shift $\Delta I(s)$ after ablating activations at idiom span, for contexts $s\in\{s_{a},s_{f},s_{l}\}$. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$, $\Delta L(s_{l})$, $\Delta F(s_{f})$ with 95% confidence intervals. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant difference between $\Delta F(s_{a})$ and the others (paired $t$-test, $p<0.05$). The difference at * marked layer is larger than the average difference across all layers.
Results Figure 3(a) shows by how much the model’s interpretation of the idiom changes when an MLP sublayer is knocked out. We observe substantial drops of $\Delta F(s_{a})$ when knocking out early MLP sublayers (0-2): from $-0.57$ to $-0.25$. At the same time, we see an increase in the probability of the literal completions: $\Delta L(s_{a})$ rises above zero, from $0.37$ to $0.18$. And these shifts are significantly different from 0-2 layers ($p<0.05$). This indicates that early layer MLPs are diagnostic for reading out the figurative meaning while potentially suppressing the literal one. For later layers ($\geq$ 4), MLP sublayer knockout does not affect the model’s interpretation, which converges to zero.
Figure 3(b) shows the effect of attention knockout, which is similar to what we observed with MLP sublayer knockout. Specifically, knocking out MHSA leads to a drop in $\Delta F(s_{a})$ and increase in $\Delta L(s_{a})$ in the early layers (0-3). The most significant difference between $\Delta F(s_{a})$ and $\Delta L(s_{a})$ is in layer 1, where $\Delta F(s_{a})$ dropped by $-0.36$ and $\Delta L(s_{a})$ increased by $0.43$ ($p<0.05$). After layer 4, knocking out MHSA has little to no effect on interpretation, as the values converge toward zero.
Together, the two interventions track the 1 idiom’s figurative meaning retrieval: Early MHSA layers (0-3) gather and bind token-wise interactions, while early MLP layers (0-2) transform those bound representations into figurative meaning, while suppressing the literal meaning. This early layer idiom meaning retrieval aligns with the findings of Haviv et al. (2022), who argue that idiomatic information is accessed in the initial layers of LLMs during inference. After layer 4, $\Delta F(s_{a})$ and $\Delta L(s_{a})$ stabilize to zero, indicating that later sublayers are not adding new information and that the representations of the idiom span tokens are no longer used by any subsequent tokens.
4.2 Disentangling idiom specific processing from general semantic processing
To validate whether early layers in MHSA and MLP components are specifically essential for idiom processing, or generally important for semantic interpretation of a sentence, we conduct additional knockout experiments on literal and figurative paraphrases ($s_{l},\,s_{f}$). If general compositional semantic processing does not need to rely on the idiom retrieval components, we would expect that knocking out them when processing non-ambiguous sentences (i.e., paraphrases) would not disrupt the inference.
ResultsFigure 3(c) shows that the drop of knocking out early MLP sublayers is significantly smaller in layers 1 and 2 for the two paraphrases $\mathrm{min}$($\Delta L(s_{l})$, $\Delta F(s_{f})$)$=-0.11$ than for the idiomatic expression, $\mathrm{min}$($\Delta F(s_{a})$)$=-0.36$ ($p<0.05$). While, at layer 0, the knockout effect is large for all expressions, as in this case, the MLP block is processing the semantics of the sentence.
Figure 3(d) shows a similar effect for the early attention layers (0-2) where literal paraphrases are not strongly affected by knockout; $\mathrm{min}$($\Delta L(s_{l})$, $\Delta F(s_{f})$)$=-0.08$, $\mathrm{min}$($\Delta F(s_{a})$)$=-0.36$ ($p<0.05$). Taken together, these results reveal that early MLP and MHSA sublayers are essential for the model’s retrieval of an idiom’s figurative meaning, and that this process is notably different for idiomatic phrases from general semantic processing. Knocking out the early layers preserves the general semantic comprehension, but disrupts its ability to interpret idiom expressions. This dissociation implies that the model activates a specialized component that is necessary for idiom understanding, yet remains inactive during non-ambiguous literal semantic processing.
4.3 Identifying attention heads specific to retrieving figurative meanings
Next, we investigated which attention heads are specialized for retrieving an idiom’s figurative meaning, while suppressing its literal compositional interpretation. We knock out attention heads iteratively at the idiom span tokens and measure $\Delta F(s_{a})$, $\Delta L(s_{a})$ to identify the attention heads most impactful for interpretation.
(a) $\Delta F(s_{a})$
(b) $\Delta L(s_{a})$
Figure 4: Heatmaps of the (a) $\Delta F(s_{a})$ (b) $\Delta L(s_{a})$ when ablating individual attention heads at the idiom span. Idiomatic heads: Heads those are crucial for retrieving figurative meaning of idiom; $-\Delta F(s_{a})$ and $+\Delta L(s_{a})$.
If a head contributes to retrieving figurative meanings from idioms, we expect that knocking it out would cause the model to lose confidence in the figurative meaning of $s_{a}$, while boosting its literal interpretation. To operationalize this, we define idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ as heads with a large negative $\Delta F(s_{a})$ and positive $\Delta L(s_{a})$. Figure 4 shows that there is indeed a subset of attention heads that simultaneously exhibits large $-\Delta F(s_{a})$ (Figure 4(a)) and $+\Delta L(s_{a})$ (Figure 4(b)). To identify which of these heads are consistently important across instances of $s_{a}$, we performed a nonparametric bootstrap with $B=1000$ resamples to identify, estimating 95% confidence intervals for each head’s contribution to the interpretation. We then rank all heads by, on the one hand, $-\Delta F(s_{a})$, and, on the other hand, $+\Delta L(s_{a})$, and select the top 20 heads highly ranked on both orders, obtaining our list of 20 idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ (selected attention heads listed in Appendix Table 1). Further analyses show that these heads process literal contexts differently compared both to random heads, and to heads that instead focus on compositional semantic processing $\mathcal{H}_{\mathrm{sem}}$ (Appendix B).
4.4 Causal role of MLP sublayers and idiomatic heads in figurative reading
To validate that the early MLP sublayers and the selected idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ are critical for the figurative reading, we patch their activations on $s_{a}$ into the same components when processing $s_{l}$. As a control, we patch randomly chosen MLP sublayers and heads. We quantify the effect on interpretation by $\Delta F(s_{l}\hookleftarrow s_{a})$ and $\Delta L(s_{l}\hookleftarrow s_{a})$.
ResultsPatching these components boosts the figurative interpretation, as quantified by $\Delta F(s_{l}\hookleftarrow s_{a})$ ($M=0.51,\,SD=0.59$) while suppressing the literal interpretation, $\Delta L(s_{l}\hookleftarrow s_{a})$ ($M=-0.92,\,SD=0.59$). In contrast, patching random components barely affects the figurative interpretation ($\Delta F(s_{l}\hookleftarrow s_{a})$: $M=0.07,\,SD=0.75$) even while suppressing the literal interpretation ($\Delta L(s_{l}\hookleftarrow s_{a})$: $M=-1.83,\,SD=1.39$). Effects on the experiment and the control are significantly different ($p<0.05$). These contrasting effects suggest that literal interpretation depends on scattered non-specialized heads, so perturbing random heads disrupts $\Delta L(s_{l})$. Conversely, idiomatic heads are specialized to enhancing the figurative meaning while suppressing literal interpretations, so that replacing their activations injects a figurative interpretation and amplifies $\Delta F(s_{l})$, while dropping $\Delta L(s_{l})$.
5 Tracing the interpretation flow
5.1 Locating interpretation representation in token position
After the retrieval process, the model encodes both literal and figurative meanings in its residual stream. We next investigate how these two representations are weighted or integrated in the final prediction. To locate the representation of the meanings, we calculate the mutual nearest-neighbor kernel alignment Huh et al. (2024); Cho et al. (2024) between sentence embeddings of meaning parts from $s_{f}$ and $s_{l}$ (i.e., exclude the ‘X (would)’ and ‘because X was so/a/the’) encoded by BGE M3 Chen et al. (2024) and hidden states from various token positions of idiom sentence $s_{a}$. This metric quantifies the similarity between representations, with higher values indicating greater semantic similarity.
(a) Last idiom token
(b) Subsequent token
(c) After subsequent token
(d) Last token of sentence
Figure 5: Kernel alignment between hidden states ($\mathrm{x}$) extracted from four different token positions of $s_{a}$ and semantic embeddings of paraphrases ($s_{f},\,s_{l}$).
ResultsWe plot the kernel alignment at four token positions in Figure 5. We observe that at the subsequent token (Fig. 5(b)), the kernel alignment between the idiom’s hidden states and its figurative meaning surpasses that of the literal meaning after layer 4. This suggests that by this point, the model has begun to favor the figurative interpretation ( 2 selective interpretation step) after retrieving figurative meaning.
[…]
Figure 6: Conceptual description of the activation patching experiments for tracing information flow (L = literal interpretation; F = figurative interpretation).
[…]
Figure 7: Layer-wise interpretation shift after patching in activations from $s_{f},\,s_{l},\,s_{a}^{*}$. The red asterisk (*) marks layers where the difference between $\Delta L$ and $\Delta F$ is larger than the average difference across all layers; these are statistically significant differences according to a paired $t$-test, $p<0.05$.
5.2 Analyzing layer-wise competing interpretation flow through activation patching
To dissociate the computational pathways of processing the figurative and literal interpretations, we employ activation patching at the subsequent token (i.e., ‘because’) across layers, the point where the figurative interpretation is most strongly encoded. We generate three conditions by pairing a source – a figurative context, a literal context, or a different idiom context ($s_{f},\,s_{l},\,s_{a}^{*}$), with a target sentence ($s_{a}$). Then, we patch the layer-wise residual activations at the ‘because’ position from the source condition into the target, see Figure 6 for illustration of expected effect. For each layer $\ell$, we measure the patching causal effect by $\Delta F$ and $\Delta L$. Figure 7 summarizes the shift of interpretation after patching. We additionally patch idiom representations into paraphrased sentences (Appendix 9), which support pathways of interpretations that we find.
ResultsPatching in figurative activations $s_{f}$ into $s_{a}$ removes the intermediate path literal flow but leaves the direct route intact (Figure 6(a)). Figure 7(a) shows that this slightly increases the $\Delta F(s_{a}\hookleftarrow s_{f})$ by up to $0.01$ for patching in mid-layers 4-10, but doesn’t affect the $\Delta L(s_{a}\hookleftarrow s_{f})$, which remains zero. This indicates that for the mid-layers at the ‘because’ token position, the model’s residual stream is the predominant channel for the figurative interpretation ( 3-1 figurative path). That is, we argue that there’s a different path for the literal interpretation ( 3-2 compositional semantic direct path) to convey its information to the final prediction, directly bypassing the intermediate path. This claim is supported by an attention-edge knockout experiment, where we prevent the last token from attending to the idiom tokens, and observe a drop in $\Delta L(s_{a})$ at layer 11 and 13 while $\Delta F(s_{a})$ remains near zero (see Appendix D).
We next patch the intermediate path by activations from the literal paraphrase $s_{l}$. This removes the intermediate path figurative flow (Figure 6(b)). Figure 7(b) shows a substantial increase in the $\Delta L(s_{a}\hookleftarrow s_{l})$, by $0.22$ in mid-layers (5-10). At the same time, $\Delta F(s_{a}\hookleftarrow s_{l})$ drops sharply at intermediate layers, which indicates that the figurative interpretation is severely disrupted.
By replacing idiom A’s intermediate representation with that of idiom B, we remove any information specific to idiom A that would have been transmitted through the intermediate pathway. Figure 7(c) shows that there is a drop of $\Delta F(s_{a}\hookleftarrow s_{a}^{*})$ by $-0.09$ in layer 7. On the other hand, the probability of the literal interpretation $\Delta L(s_{a}\hookleftarrow s_{a}^{*})$ does not change significantly. This implies that the literal path is intact, and there’s a bit of noise by adding the unrelated activations which cause a slight drop of $\Delta L(s_{a}\hookleftarrow s_{a}^{*})$ compared to $\Delta L(s_{a}\hookleftarrow s_{l})$. The magnitude of the drop in $\Delta F(s_{a}\hookleftarrow s_{a}^{*})$ closely matches the decrease after patching in the literal paraphrase $\Delta F(s_{a}\hookleftarrow s_{l})$, demonstrating that the figurative interpretation is predominantly routed through the intermediate pathway. In Appendix D, we further demonstrate that ablating the attention edge from the idiom span to its subsequent token causes $\Delta L(s_{a})$ to increase while $\Delta F(s_{a})$ decreases. This provides additional evidence that the figurative meaning is predominantly conveyed through this intermediate pathway. Meanwhile, when we patch activations from idiom B into the intermediate pathway of idiom A, we expect the intermediate representations to be overwritten by those of idiom B (Figure 6(c)). This would lead to increased prediction probabilities for both the figurative and literal interpretations associated with idiom B. As shown in Figure 7(d), both $\Delta F^{B}(s_{a}\hookleftarrow s_{a}^{*})$ and $\Delta L^{B}(s_{a}\hookleftarrow s_{a}^{*})$ increase, indicating that information supporting both interpretations is transmitted through the intermediate path. However, the magnitudes of these increases are not equal, suggesting that the pathway preferentially amplifies figurative interpretation over the other.
6 Conclusion
We present multiple steps of idiom processing in a transformer model (Figure 1), and identify distinct components and flows involved in the interpretation of figurative meaning. By isolating idiomatic expressions from the context, we gain insight into which components of the model are most sensitive to figurative interpretation compared to literal one. However, contextual information plays a crucial role in interpreting idioms, as it helps determine whether a given expression should be understood in its literal or figurative sense Holsinger (2013). Future research would explore how the model’s processing changes when disambiguating context is provided. By inserting disambiguating cues both before and after the idiom, we could examine the interactions between contextual clarification and idiom comprehension.
Moreover, an important next step is to extend our analysis to a larger number of idioms and develop a diagnostic framework that predicts, from the model’s internal activations alone, whether an expression’s figurative meaning is known. Concretely, one could observe the engagement of idiom-specialized attention heads and the activation of retrieval and binding pathways for every idiom in the set. If a given idiom fails to activate its dedicated heads, this would point to a breakdown in retrieval or binding. By looking at these activation patterns, we could quantify the model’s competence with respect to the idiom comprehension. Such an approach would not only deepen our understanding of how figurative meanings are encoded in transformer representations but also provide a useful guidance for pinpointing specific processing failures in idiom comprehension.
Limitations
While causal tracing methods have been widely used in recent work Geva et al. (2023); Dar et al. (2022); Meng et al. (2022); Heimersheim and Nanda ; Nanda et al. (2023), they only approximate the actual information stored in activations. Moreover, knocking out or replacing the activations can lead the model to out-of-distribution behaviors and cast doubt on the robustness of any interpretability claim. Our study is limited to LLama3.2-1B-base model, which is much smaller than the current state-of-the-art transformers, scaling these analyses to larger models remains an open challenge. Also, we simplify idiom processing by excluding context, even though context critically shapes idiomatic meaning. Altogether, these factors limit the generalizability of our results.
Ethics statement
One of the intended uses of the LLama-3.3-70B-instruct model is content generation, which aligns with our use of the model for generating the candidates for next token prediction task Grattafiori et al. (2024). The candidates generated by the language model may include biased or sensitive attributes (e.g., race, minority status), which reflects stereotypes that the language model already has (See the Appendix Table 3 for examples).
Acknowledgements
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Starting Grant “Individualized Interaction in Discourse”, grant agreement No. 948878). Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102.
Appendix A Selected top-20 attention heads
We further extracted semantic heads $\mathcal{H}_{\mathrm{sem}}$, i.e., heads that show a positive $\Delta F(s_{a})$ and a negative $\Delta L(s_{a})$, which are processing ambiguous sentence in a literal way. Ablating them suppresses literal processing and restores figurative understanding. For comparison, we also randomly selected heads that are neither in $\mathcal{H}_{\mathrm{idiom}}$ nor $\mathcal{H}_{\mathrm{sem}}$.
Type top-$20$ attention heads (layer, head)
Idiomatic (0, 4), (0, 28), (1, 5), (2, 9), (0, 30), (0, 8), (0, 19), (2, 2), (1, 25), (1, 18), (4, 31), (1, 31), (0, 12), (0, 3), (2, 28), (2, 17), (2, 8), (2, 30), (1, 9), (4, 24)
Semantic (1, 3), (3, 28), (3, 30), (0, 10), (7, 16), (1, 20), (2, 4), (3, 12), (3, 14), (6, 27), (2, 19), (10, 11), (4, 26), (0, 18), (6, 31), (9, 22), (2, 5), (0, 9), (1, 19), (0, 11)
Random (9, 20), (10, 30), (2, 11), (2, 23), (1, 1), (9, 23), (13, 24), (8, 31), (3, 8), (13, 9), (10, 1), (5, 15), (7, 14), (7, 24), (4, 5), (14, 2), (9, 17), (4, 0), (14, 19), (6, 9)
Table 1: Top-$20$ attention head sets for each type.
Appendix B Activation divergence analysis between idiomatic and literal contexts
To examine how idiomatic heads $\mathcal{H}_{\mathrm{idiom}}$ process idioma differently compared to literal contexts, we focus on their activations over the idiom token span. Let =$\mathcal{I}=\{i_{1},\dots,i_{K}\}\subseteq\{1,\dots,T\}$ be the indices of the idiom span tokens. Note that the literal paraphrase $s_{l}$ preserves the same idiom token span, hence $|\mathcal{I}|$ is identical for both $s_{a}$ and $s_{l}$. For each head $h\in\mathcal{H}_{\mathrm{idiom}}$, we compute the averaged activation from idiom span value vectors:
$v_{a}\;=\;\frac{1}{\mathcal{I}}\sum_{i\in\mathcal{I}}\alpha^{(a)}_{h,i}\,V_{i},\;v_{l}\;=\;\frac{1}{\mathcal{I}}\sum_{i\in\mathcal{I}}\alpha^{(l)}_{h,i}\,V_{i},$
where $\alpha^{(o)}_{h,i}$ and $\alpha^{(l)}_{h,i}$ are the attention weights on token $i$ in $s_{a}$ and $s_{l}$, respectively, and $V_{i}$ is the corresponding value projection. These mean vectors capture the direction in the head’s value subspace over the idiom span under each context. We then measure their alignment via cosine similarity:
$\cos(v_{a},\,v_{l})=\frac{v_{a}\cdot v_{l}}{\|v_{a}\|\,\|v_{l}\|}.$
A low cosine similarity between $v_{a}$ and $v_{l}$ means that the head’s representational direction shifts substantially when processing the idiomatic sentence $s_{a}$ compared to its literal paraphrase $s_{l}$, revealing its specialization in processing ambiguous sentence and extracting its figurative meaning. In contrast, a high cosine similarity indicates that the head maintains a similar representation across both contexts, reflecting a more general semantic role. For comparison, we also compute cosine similarities for the semantic heads $\mathcal{H}_{\mathrm{sem}}$ and for a random set of heads $\mathcal{H}_{\mathrm{rand}}$.
For each sentence, we compute the cosine similarity $\cos(v_{a},v_{l})$ for every head in a given set of 20 attention heads, and then average these 20 values to obtain a single mean similarity (Figure 8). The cosine similarity across the dataset is shown in Figure 8. The $cos\_random$ ($M=0.87,\,SD=0.07$), $cos\_sem$ ($M=0.80,\,SD=0.09$) is significantly larger than $cos\_idiom$ ($M=0.76,\,SD=0.12$), ($p<0.05$).
Figure 8: Cosine similarities between the attention activations on idiomatic sentences ($s_{a}$) and their literal paraphrase ($s_{l}$) for the top-20 idiomatic heads, semantic heads, and random heads.
By confirming that the heads identified as specialized in retrieving the idioms’ figurative meaning indeed have divergent attention patterns when processing idioms compared to literal paraphrases, we strengthen the evidence that these heads are responsible for the unique processing required to interpret idiomatic expressions correctly.
Appendix C Layer-wise interpretation flow through activation patching
We additionally patch idiom’s because token activation into paraphrased sentences’ because token position. The conceptual details of the experiment is described as follow.
[…]
Figure 9: Description of the activation patching experiment for identifying interpretation flow. L represents literal interpretation; F represents figurative interpretation. (a) By replacing ‘because’ token $s_{f}$ activation with $s_{a}$, we expect literal interpretation to be added in the intermediate channel while keeping the direct route flow to be activated with naive figurative meaning. (b) By replacing ‘because’ token $s_{l}$ activation with $s_{a}$, we expect figurative interpretation to be added via intermediate path, while keeping both paths for literal interpretation.
[…]
Figure 10: Layer-wise interpretation shift $\Delta I(s_{\mathrm{x}}\hookleftarrow s_{a}),\,\mathrm{x}\in\{f,l\}$ after activation patching on the ‘because’ token across three patch directions. Y-axis: Mean values of $\Delta L(s_{\mathrm{x}}\hookleftarrow s_{a})$, $\Delta F(s_{\mathrm{x}}\hookleftarrow s_{a})$. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant shift between literal ($C_{l}$) and figurative ($C_{f}$) candidate probability (paired $t$-test, $p<0.05$).
As expected, by adding idiom’s interpretation via intermediate channel, since it’s adding literal interpretation, there’s a increase in $\Delta L(x_{f}\hookleftarrow s_{a})$ by $0.1$, whereas $\Delta F(x_{f}\hookleftarrow s_{a})$ drops by $-0.07$ across layers (Figure 10(a)). Whereas, when adding the idiom’s figurative interpretation into intermediate path, there’s a big increase in $\Delta F(s_{l}\hookleftarrow s_{a})$ by $0.26$ at layer 7 and decrease in $\Delta L(s_{l}\hookleftarrow s_{a})$ by $-0.17$. This implies that when there’s idiomatic representation passing the intermediate path, to make final prediction as close to figurative one, the model tries to amplify figurative meaning that is passing the intermediate path so that the total amount of activations for figurative interpretation to be same, which should be more than its literal one.
That is, injecting the idiom’s literal activation into the intermediate channel increases $\Delta L(x_{f}\hookleftarrow s_{a})$ by up to $0.1$ in the mid-layers (4-9). Conversely, injecting figurative activation raises $\Delta F(s_{l}\hookleftarrow s_{a})$ by $0.26$. This asymmetric increase indicates that when an idiomatic representation passes the intermediate path, the model actively amplifies figurative interpretation so it dominates the final prediction.
Appendix D Analyzing edge-level intermediate channel through attention knockout
To trace the flow of information, we perform fine-grained intervention on the MHSA sublayers, since they mediate communication between token positions, therefore, any critical information must pass through them Geva et al. (2023).
To locate the intermediate and direct path for each interpretation, we focus on the attention scores between idiom to subsequent token (i.e., because) for the intermediate path, and the idiom to last token of the sentence for the direct path findings. We investigate how this interpretation of flow is done internally by knockout the attention edges between target tokens. When critical attention edges for a certain interpretation of the idiom are blocked, it will result in severe degradation in probabilities of certain candidates. Therefore, we test whether critical information propagates between two hidden representations at a specific layer, by zeroing-out all the attention edges between them.
Formally, let $i,j\in\{1,\dots,T\}$ with $j\leq i$ be two positions in the sequence. To block token $i$ (query) from attending to token $j$ (key) at layer $\ell<L$, we overwrite the corresponding entries in the causal self-attention mask $M$ at layer $\ell+1$ across all heads $h\in\{1,\dots,H\}$:
$M_{i,\,j}^{\ell+1,\,h}\;=\;-\,\infty\quad\forall\,h\in\{1,\dots,H\}.$
(a) Idiom $\rightarrow$ Subsequent
(b) Idiom $\rightarrow$ Last
Figure 11: Layer-wise interpretation shift $\Delta I(s_{a})$ after knock out the attention edges between tokens. Y-axis: Mean values of $\Delta L(s_{a})$, $\Delta F(s_{a})$. X-axis: Layers. Gray dashed line: $\Delta I=0$ (no effect). Red asterisk (*): Significant shift between literal ($C_{l}$) and figurative ($C_{f}$) candidate cumulative probabilities (paired $t$-test, $p<0.05$).
This restricts the query position from obtaining information from the key position, at that particular layer ($\ell+1$). For each layer, we block attention edges between (a) idiom span tokens $\rightarrow$ subsequent token (b) idiom span tokens $\rightarrow$ last token. If the specific path is important for a specific interpretation, then there should be significant drop of $I(s_{a})$ value. Figure 11 shows the result.
In Figure 11(a) at layer 5, it shows that there’s a drop in the $\Delta F(s_{a})$ by $-0.02$, while increase in $\Delta L(s_{a})$, by $0.02$. This shows that ablating attention from the idiom span to ‘because’ token disrupts the indirect path that would do disambiguation process where it suppresses the literal one. Without the idiom attending to the ‘because’ token, the model can no longer effectively disambiguate figurative meaning from literal meaning.
At layer 11 and 13, as in Figure 11(b), $\Delta L(s_{a})$ declines while $\Delta F(s_{a})$ rises. This divergence indicates that ablating the direct idiom to last token pathway prevents the model from passing the literal interpretation.
Ambiguous sentence Literal paraphrase Figurative paraphrase Literal candidates Figurative candidates
They will bend over backwards because they are so They will arch spine backwards because they are so They will make extra efforts because they are so flexible, used, strong, weak, relaxed, tight, uncomfortable, tall, stiff, short, scared, full, comfortable, tense, small, thin, angry, over, inf, surprised grateful, eager, proud, passionate, close, motivated, committed, keen, desperate, enthusiastic, invested, confident, glad, well, interested, sure, apprec, focused, attracted, dedicated
He bit off more than he can chew because he was so He took more food than he can swallow because he was so He took on more than he could handle because he was so hungry, greedy, happy, very, nervous, fam, attracted, r, thirsty, poor, star, pleased, starving, sad, delighted, hung, eng, drunk, gl, tempted eager, confident, anxious, desperate, optimistic, enthusiastic, passionate, sure, determined, ambitious, focused, driven, busy, good, full, young, keen, convinced, smart, strong
He would blow his own horn because he was a He would blow the musical instrument because he was a He would praise himself because he was a professional, musician, fan, p, wind, skilled, shepherd, trumpet, member, trump, pro, fl, virt, jazz, boy, bag, human, brass, flute, musical good, great, man, self, genius, proud, winner, god, true, hero, hard, legend, unique, better, brilliant, successful, clever, narciss, smart, perfection
It was out in left field because it was a It sat far in baseball’s area because it was a It was completely unrealistic because it was a baseball, league, minor, sport, strong, few, popular, difficult, download, home, football, smaller, tough, sports, right, significant, basketball, pitcher, deep, stadium fantasy, dream, very, completely, huge, one, product, total, movie, story, complete, perfect, romantic, fairy, totally, cartoon, fictional, massive, two, film
They were up the creek because they were a They were near the stream because they were a They were in trouble because they were a group, family, fishing, nom, part, hunting, water, tribe, pair, thirsty, party, people, river, fish, stream, traveling, farming, curious, type, peaceful new, long, threat, small, minority, bunch, few, mixed, very, day, single, large, bad, man, young, mess, tiny, mix, poor, relatively
Table 2: Examples of generated data.
Ambiguous sentence Literal paraphrase Figurative paraphrase Literal candidates Figurative candidates
They left him out in the cold because he was a They left him outside in the frost because he was a They abandoned him without support because he was a bad, stranger, little, bit, thief, trouble, poor, drunk, witch, dirty, dog, rebel, beg, danger, trait, s, nuisance, tiny, he, naughty child, foreign, threat, boy, political, baby, black, non, reminder, disabled, male, burden, cripp, product, minority, son, member, difficult, minor, different
They let him off the hook because he was a They removed him from fishing tackle because he was a They freed him from responsibility because he was a threat, bad, convicted, little, bit, danger, sexual, fish, ped, white, criminal, racist, bully, black, terrorist, thief, political, jerk, rap, big minor, child, foreign, slave, kid, good, mad, juvenile, victim, boy, stranger, young, youth, mere, student, fool, first, teenager, prisoner, friend
She was looking for a needle in a haystack because she was a She was searching for a neddle within a straw because she was a She was facing an impossible search because she was a needle, craft, seam, straw, tiny, hay, detective, master, haystack, witch, farmer, pro, camel, professional, busy, crazy, giant, chicken, cow, neat woman, victim, single, young, black, female, girl, slave, mother, stranger, new, first, non, private, white, novice, minority, prisoner, foreign, mom
Table 3: Biased data examples reflecting stereotypes embedded in the language model.